The ΔG for the freezing of H2O(l) at -10 ∘C is-210 J/mol and the heat of fusion of ice at this temperature is 5610 J/mol .

Find the entropy change of the universe when 3 mol of water freezes at -10 ∘C.
Answer #1

Calculation of delta G and delta H for 3 mol water

Calculation of T in K

T (K) = - 10 + 273.15 = 263.15 K

Delta G = 3 mol water x -210 J/mol=-630 J

Delta H = 3 mol waer x 5610 J/mol = 16830 J

Lets find delta S by using following equation.

Delta G = Delta H – Tdelta S

T delta S = Delta H – Delta G

Delta S = [Delta H – Delta G]/T

= {[-630 - ( 16830)] /263.15 }

=-66.35 J /K

Delta S = -66.35 J/K